GROTHENDIECK CATEGORIES OF ENRICHED FUNCTORS

HASSAN AL HWAEER AND GRIGORY GARKUSHA

Abstract. It is shown that the category of enriched functors [C ,V ] is Grothendieck whenever V is a closed symmetric monoidal Grothendieck category and C is a category enriched over V . Localizations in [C ,V ] associated to collections of objects of C are studied. Also, the category of chain complexes of generalized modules Ch(C_{R}) is shown to be identified with the Grothendieck category of enriched functors [modR,Ch(ModR)] over a commutative ring R, where the category of finitely presented R-modules modR is enriched over the closed symmetric monoidal Grothendieck category Ch(ModR) as complexes concentrated in zeroth degree. As an application, it is proved that Ch(C_{R}) is a closed symmetric monoidal Grothendieck model category with explicit formulas for tensor product and internal Hom-objects. Furthermore, the class of unital algebraic almost stable homotopy categories generalizing unital algebraic stable homotopy categories of Hovey–Palmieri–Strickland [14] is introduced. It is shown that the derived category of generalized modules D(C_{R}) over commutative rings is a unital algebraic almost stable homotopy category which is not an algebraic stable homotopy category.

1. Introduction

In the present paper we study categories of enriched functors

[C ,V ],

where V is a closed symmetric monoidal Grothendieck category and C is a category enriched over V (i.e. a V -category). The main result here states that the category [C ,V ] is Grothendieck with an explicit collection of generators. Namely, the following theorem is true. Theorem. Let V be a closed symmetric monoidal Grothendieck category with a set of generators {g_{i}}_{I}. If C is a small V -category, then the category of enriched functors [C ,V ] is a Grothendieck V -category with the set of generators {V (c,−)⊘g_{i} | c ∈ ObC ,i ∈ I}. Moreover, if C is a small symmetric monoidal V -category, then [C ,V ] is closed symmetric monoidal with explicit formulas for monoidal product and internal Hom-object.

Taking into account this theorem, we refer to [C ,V ] as a Grothendieck category of enriched functors. The usual Grothendieck category of additive functors

(B,Ab)

from a pre-additive category B to abelian groups Abis recovered from the preceding theorem in the case when V = Ab (B is a V -category). Further examples on how the category [C ,V ] recovers some Grothendieck categories are given in Section 4.

The following result is an extension to enriched categories of similar results of [6, 9].

Date: 15 September 2014. 2010 Mathematics Subject Classification. 18E15, 18E30, 18G55. Key words and phrases. Grothendieck categories, enriched categories, model categories. The first author was supported by the Ministry of Higher Education and Mathematics Department of Kufa Uni-

versity, Iraq.

Theorem. Suppose V is a closed symmetric monoidal Grothendieck category. Let C be a V - category and let P consist of a collection of objects of C. Let S_{P} = {G ∈ [C ,V ]| G(p) = 0 for all p ∈ P}. Then S_{P} is a localizing subcategory of [C ,V ]and [P,V ] is equivalent to the quotient category [C ,V ]/S_{P} .

We apply Grothendieck categories of enriched functors to study homological algebra for generalized modules. The category of generalized modules

C_{R}= (modR,Ab)

consists of the additive functors from the category of finitely presented R-modules, modR, to the category of abelian groups, Ab. Its morphisms are the natural transformations of functors. It is called the category of generalized R-modules for the reason that there is a fully faithful, right exact functor

M ↦ − ⊗_{R}M

from the category of all R-modules to C_{R}.

The category C_{R} has a number of remarkable properties which led to powerful applications in ring and module theory and representation theory (see, e.g., the books by Prest [20, 21]). The category C_{R} also provides a natural connecting langauge between algebra and model theory of modules [12, 20, 21].

The following theorem states that the category Ch(C_{R}) of chain complexes of C_{R} over a commutative ring can be regarded as a Grothendieck category of enriched functors.

Theorem. Suppose R is a commutative ring. Then the category of chain complexes of generalized R-modules Ch(C_{R}) can naturally be identified with the Grothendieck category of enriched functors [modR,Ch(ModR)], where the category of finitely presented modules modR is naturally enriched over Ch(ModR) as complexes concentrated in zeroth degree.

A Grothendieck category of enriched functors [C ,V ] can also contain a homotopy information whenever V is a reasonable model category in the sense of Quillen [22]. As an application of the preceding theorem, we show the following

Theorem. Let R be a commutative ring, then Ch(C_{R}) is a left and right proper closed symmetric monoidal V -model category, where V = Ch(ModR). The tensor product of two complexes F_{•},G_{•} ∈ Ch(C_{R}) is given by

F_{•}⊙ G_{•} = ∫ ^{(M,N)∈modR⊗modR}F_{•}(M) ⊗_{R}G_{•}(N) ⊗_{R}Hom_{R}(M ⊗_{R}N,−).

Here Hom_{R}(M ⊗_{R}N,−) is regarded as a complex concentrated in zeroth degree. The internal Hom-object is defined as

\underline{Hom}(F_{•},G_{•})(M) = ∫_{N∈modR} \underline{Hom}_{Ch(ModR)}(F_{•}(N),G_{•}(M ⊗_{R}N)).

There are various ways to construct closed symmetric monoidal structures on the derived category of a reasonable closed symmetric abelian category (see [13, 25] for some examples). For the case of the derived category D(C_{R}) of generalized modules with R a commutative ring we apply the preceding theorem as well as some facts for compactly generated triangulated categories to establish the following

Theorem. Let R be a commutative ring. Then the derived category D(C_{R}) of the Grothendieck category C_{R} is a compactly generated triangulated closed symmetric monoidal category, where the above formulas yield the derived tensor product F_{•}⊙^{L}G_{•} and derived internal Hom-object R\underline{Hom}(F_{•},G_{•}). The compact objects of D(C_{R}) are the complexes isomorphic to bounded complexes of coherent functors.

In the classical stable homotopy theory (see, for example, Hovey–Palmieri–Strickland [14]) the category of compact objects of a stable homotopy category possesses a duality, which in some cases is also known as the Spanier–Whitehead Duality. In order to find a duality on the category of compact objects D(C_{R})^{c} of D(C_{R}), we use the Auslander–Gruson–Jensen Duality [1, 10, 12] for coherent objects cohC_{R}. In the model theory of modules, this duality corresponds to elementary duality, introduced by Prest [20, Chapter 8] and developed by Herzog in [11], for positive-primitive formulas. We show that the Auslander–Gruson–Jensen Duality makes sense for compact objects of D(C_{R}). More precisely, the following result is true.

Theorem (Auslander–Gruson–Jensen Duality for compact objects). Let D(C_{R})^{c} be the full triangulated subcategory of D(C_{R}) of compact objects. Then there is a duality

D: (D(C_{R})^{c})^{op} → D(C_{R})^{c}

that takes a compact object C_{•} to DC_{•} := R\underline{Hom}(C_{•},− ⊗_{R}R).

Basing on the above results for D(C_{R}), we introduce the class of unital algebraic almost stable homotopy categories. These essentially the same with unital algebraic stable homotopy categories in the sense of Hovey–Palmieri–Strickland [14] except that the compact objects do not have to be strongly dualizable, but must have a duality. We finish the paper by proving the following

Theorem. Let R be a commutative ring. Then D(C_{R}) is a unital algebraic almost stable homotopy category, which is not an algebraic stable homotopy category in the sense of Hovey– Palmieri–Strickland.

2. Enriched Category Theory

In this section we collect basic facts about enriched categories we shall need later. We refer the reader to [2, 18] for details. Throughout this paper (V ,⊗,\underline{Hom},e) is a closed symmetric monoidal category with monoidal product ⊗, internal Hom-object \underline{Hom}and monoidal unit e. We sometimes write [a,b] to denote \underline{Hom}(a,b), where a,b ∈ ObV . We have structure isomorphisms in V

a_{abc}: (a ⊗ b) ⊗ c → a ⊗ (b ⊗ c), l_{a} : e ⊗ a → a, r_{a} : a ⊗ e → a

with a,b,c ∈ ObV .

Definition 2.1. A V -category C, or a category enriched over V , consists of the following data: (1) a class Ob(C)of objects;

(2) for every pair a,b ∈ Ob(C)of objects, an object V_{C}(a,b) of V ;

(3) for every triple a,b,c ∈ Ob(C)of objects, a composition morphism in V ,

c_{abc}: V_{C}(a,b) ⊗ V_{C}(b,c) → V_{C}(a,c);

(4) for every object a ∈ C, a unit morphism u_{a} : e → V_{C}(a,a) in V . These data must satisfy the following conditions: ⋄ given objects a,b,c,d ∈ C, diagram (1) below is commutative (associativity ax-

iom);

⋄ given objects a,b ∈ C, diagram (2) below is commutative (unit axiom).

\xymatrix{(V_{C}(a,b) ⊗ V_{C}(b,c)) ⊗ V_{C}(c,d) \ar[r]^{c_{abc}⊗1} \ar[d]_{a_{V_{C}(a,b)V_{C}(b,c)V_{C}(c,d)}} & V_{C}(a,c) ⊗ V_{C}(c,d) \ar[dd]^{c_{acd}} \\ V_{C}(a,b) ⊗ (V_{C}(b,c) ⊗ V_{C}(c,d)) \ar[d]_{1⊗c_{bcd}} \\ V_{C}(a,b) ⊗ V_{C}(b,d) \ar[r]_{c_{abd}} & V_{C}(a,d)}  (1)

\xymatrix{e ⊗ V_{C}(a,b) \ar[r]^{l_{V_{C}(a,b)}} \ar[d]_{u_{a}⊗1} & V_{C}(a,b) \ar[d]^{1_{V_{C}(a,b)}} & V_{C}(a,b) ⊗ e \ar[l]_{r_{V_{C}(a,b)}} \ar[d]^{1⊗u_{b}} \\ V_{C}(a,a) ⊗ V_{C}(a,b) \ar[r]^{c_{aab}} & V_{C}(a,b) & V_{C}(a,b) ⊗ V_{C}(b,b) \ar[l]_{c_{abb}}}  (2)

When ObC is a set, the V -category C is called a small V -category.

Definition 2.2. Given V -categories A ,B, a V -functor or an enriched functor F : A → B consists in giving:

(1) for every object a ∈ A, an object F (a) ∈ B;

(2) for every pair a,b ∈ A of objects, a morphism in V ,

F_{ab}: V_{A} (a,b) → V_{B} (F (a),F (b))

in such a way that the following axioms hold: ⋄ for all objects a,a′,a′′∈ A, diagram (3) below commutes (composition axiom); ⋄ for every object a ∈ A, diagram (4) below commutes (unit axiom).

\xymatrix{V_{A} (a,a′) ⊗ V_{A} (a′,a′′) \ar[r]^{c_{aa′a′′}} \ar[d]_{F_{aa′}⊗F_{a′a′′}} & V_{A} (a,a′′) \ar[d]^{F_{aa′′}} \\ V_{B} (Fa,Fa′) ⊗ V_{B} (Fa′,Fa′′) \ar[r]_{c_{Fa,Fa′,Fa′′}} & V_{B} (Fa,Fa′′)}  (3)

\xymatrix{e \ar[r]^{u_{a}} \ar[dr]_{u_{Fa}} & V_{A} (a,a) \ar[d]^{F_{aa}} \\  & V_{B} (Fa,Fa)}  (4)

Definition 2.3. Let A ,B be two V -categories and F,G: A → B two V -functors. A V -natural transformation α : F ⇒ G consists in giving, for every object a ∈ A, a morphism

α_{a} : e → V_{B} (F (a),G(a))

in V such that diagram (5) below commutes, for all objects a,a′∈ A.

\xymatrix{ & V_{A} (a,a′) \ar[dl]_{l_{V_{A} (a,a′)}^{−1}} \ar[dr]^{r_{V_{A} (a,a′)}^{−1}} \\ e ⊗ V_{A} (a,a′) \ar[d]_{α_{a}⊗G_{aa′}} &  & V_{A} (a,a′) ⊗ e \ar[d]^{F_{aa′}⊗α_{a′}} \\ V_{B} (Fa,Ga) ⊗ V_{B} (Ga,Ga′) \ar[dr]_{c_{FaGaGa′}} &  & V_{B} (Fa,Fa′) ⊗ V_{B} (Fa′,Ga′) \ar[dl]^{c_{FaFa′Ga′}} \\  & V_{B} (Fa,Ga′)} (5)

By Set we shall mean the closed symmetric monoidal category of sets. Categories in the usual sense are Set-categories (categories enriched over Set). If A is a category, let Set_{A} (a,b) denote the set of maps in A from a to b. The closed symmetric monoidal category V is a V -category due to its internal Hom-objects. Any V -category C defines a Set-category U C, also called the underlying category. Its class of objects is ObC, the morphism sets are Set_{U C}(a,b) = Set_{V}(e,V_{C}(a,b))(see [2, p. 316]).

Proposition 2.4. Let V be a symmetric monoidal closed category. If A is a V -category and F,G: A ⇉ V are V -functors, giving a V -natural transformation α : F ⇒ G is equivalent to giving a family of morphism α : F (a) → G(a) in V , for a ∈ A, in such a way that the following diagram commutes for all a,a′∈ A

\xymatrix{V_{A} (a,a′) \ar[r]^{F_{aa′}} \ar[d]_{G_{aa′}} & [F (a),F (a′)] \ar[d]^{[1,α_{a′}]} \\ [G(a),G(a′)] \ar[r]_{[α_{a},1]} & [F (a),G(a′)]}

Proof. See [2, 6.2.8]. □

Corollary 2.5. Let V be a symmetric monoidal closed category. If A is a V -category and F,G: A ⇉ V are V -functors, giving a V -natural transformation α : F ⇒ G is equivalent to giving a family of morphism α : F (a) → G(a) in V , for a ∈ A, in such a way that the following diagram commutes for all a,a′∈ A

\xymatrix{V_{A} (a,a′) ⊗ F (a) \ar[r]^{η_{F}} \ar[d]_{1⊗α_{a}} & F (a′) \ar[d]^{α_{a′}} \\ V_{A} (a,a′) ⊗ G(a) \ar[r]_{η_{G}} & G(a′),}

where η_{F},η_{G} are the morphisms corresponding to the structure morphisms F_{aa′} and G_{aa′} respectively.

Corollary 2.6. Let V be a symmetric monoidal closed category. If A is a small V -category and F,G : A ⇉ V are V -functors. Suppose α : F ⇒ G is a V -natural transformation such that each α_{a} : F (a) → G(a), a ∈ ObA, is an isomorphism in V . Then α is an isomorphism in [A ,V ].

Proof. This follows from the preceding corollary if we define α^{−1} : G ⇒ F by the collection of arrows α_{a}^{−1} , a ∈ ObA. □

Let C ,D be two V -categories. The monoidal product C ⊗ D is the V -category, where

Ob(C ⊗ D) := ObC × ObD

and

V_{C ⊗D} ((a,x),(b,y)) := V_{C}(a,b) ⊗ V_{D}(x,y), a,b ∈ C ,x,y ∈ D.

Definition 2.7. A V -category C is a right V -module if there is a V -functor act: C ⊗ V → C, denoted (c,A) ↦ c ⊘ A and a V -natural unit isomorphism r_{c} : act(c,e) → c subject to the following conditions:

(1) there are coherent natural associativity isomorphisms c ⊘ (A ⊗ B) → (c ⊘ A) ⊗ B;

(2) the isomorphisms c ⊘ (e ⊗ A) ⇉ c ⊘ A coincide.

A right V -module is closed if there is a V -functor

coact: V ^{op}⊗ C → C

such that for all A ∈ ObV , and c ∈ ObC, the V -functor act(−,A) : C → C is left V -adjoint to coact(A,−) and act(c,−) : V → C is left V -adjoint to V_{C}(c,−).

If C is a small V -category, V -functors from C to V and their V -natural transformations form the category [C ,V ] of V -functors from C to V . If V is complete, then [C ,V ] is also a V -category. We also denote this V -category by F (C), or F if no confusion can arise. The morphism V -object V_{F}(X,Y) is the end

∫_{ObC} V (X(c),Y (c)).  (6)

Note that the underlying category U F of the V -category F is [C ,V ].

Given c ∈ ObC, X ↦ X(c) defines the V -functor Ev_{c} : F → V called evaluation at c. The assignment c ↦ V_{C}(c,−) from C to F is again a V -functor C^{op} → F, called the V -Yoneda embedding. V_{C}(c,−) is a representable functor, represented by c.

Lemma 2.8 (Enriched Yoneda Lemma). Let V be a complete closed symmetric monoidal category and C a small V -category. For every V -functor X : C → V and every c ∈ ObC, there is a V -natural isomorphism X(c) ≅ V_{F} (V_{C}(c,−),X).

Lemma 2.9. If V is a bicomplete closed symmetric monoidal category and C is a small V - category, then [C ,V ] is bicomplete. (Co)limits are formed pointwise.

Proof. See [2, 6.6.17]. □

Corollary 2.10. Assume V is bicomplete, and let C be a small V -category. Then any V -functor X : C → V is V -naturally isomorphic to the coend

X ≅ ∫ ^{ObC} V_{C}(c,−) ⊘ X(c).

Proof. See [2, 6.6.18]. □

A monoidal V -category is a V -category C together with a V -functor ⋄ : C ⊗ C → C, a unit u ∈ ObC, a V -natural associativity isomorphism and two V -natural unit isomorphisms. Symmetric monoidal and closed symmetric monoidal V -categories are defined similarly.

Suppose (C ,⋄,u) is a small symmetric monoidal V -category where V is bicomplete. In [4], a closed symmetric monoidal product was constructed on the category [C ,V ] of V -functors from C to V . For X,Y ∈ Ob[C ,V ], the monoidal product X ⊙ Y ∈ Ob[C ,V ] is the coend

X ⊙ Y := ∫ ^{Ob(C ⊗C)} V_{C}(c ⋄ d,−) ⊗ (X(c) ⊗ Y (d)) : C → V .  (7)

The following theorem is due to Day [4] and plays an important role in our analysis.

Theorem 2.11 (Day [4]). Let (V ,⊗,e) be a bicomplete closed symmetric monoidal category and (C ,⋄,u) a small symmetric monoidal V -category. Then ([C ,V ],⊙,V_{C}(u,−)) is a closed symmetric monoidal category. The internal Hom-functor in [C ,V ] is given by the end

F (X,Y)(c) = V_{F} (X,Y (c ⋄ −)) = ∫_{d∈ObC} V (X(d),Y (c ⋄ d)).  (8)

The next lemma computes the tensor product of representable V -functors.

Lemma 2.12. The tensor product of representable functors is again representable. Precisely, there is a natural isomorphism

V_{C}(c,−) ⊙ V_{C}(d,−) ≅ V_{C}(c ⋄ d,−).

3. Grothendieck categories

In this section we collect basic facts about Grothendieck categories. We mostly follow Herzog [12] and Stenström [24].

Definition 3.1. A family {U_{i}}_{I} of objects of an abelian category A is a family of generators for A if for each non-zero morphism α : B → C in A there exist a morphism β : U_{i} → B for some i ∈ I, such that αβ ≠ 0.

Recall that an abelian category is cocomplete or an Ab3-category if it has arbitrary direct sums. The cocomplete abelian category C is said to be an Ab5-category if for any directed family {A_{i}}_{i∈I} if subobjects of A and for any subobject B of A, the relation

(∑_{i∈I} A_{i}) ∩ B = ∑_{i∈I}(A_{i}∩ B)

holds.

The condition Ab3 is equivalent to the existance of arbitrary direct limits. Also Ab5 is equivalent to the fact that there exist inductive limits and the inductive limits over directed families of indices are exact, i.e. if I is a directed set and

\xymatrix{0 \ar[r] & A_{i} \ar[r] & B_{i} \ar[r] & C_{i} \ar[r] & 0}

is an exact sequence for any i ∈ I, then

\xymatrix{0 \ar[r] & \underrightarrow{lim}A_{i} \ar[r] & \underrightarrow{lim}B_{i} \ar[r] & \underrightarrow{lim}C_{i} \ar[r] & 0}

is an exact sequence.

An abelian category which satisfies the condition Ab5 and which possesses a family of generators is called a Grothendieck category.

Example 3.2. Given any associative ring R with identity, the category ModR of R-modules is a Grothendieck category, where R is a generator.

Example 3.3. By [12], another example of a Grothendieck category is the category of additive functors (B,Ab) from a small preadditive category B to the category of abelian groups Ab, in which limits and colimits of functors are defined objectwise. A family of projective generators for (B,Ab) is given by the collection of representable functors {h^{B}}_{B∈ObB} . In what follows we shall also write (B,−) to denote the representable functor h^{B}, B ∈ ObB.

Below we shall need the following useful fact:

Proposition 3.4. The category of unbounded chain complexes Ch(A) of a Grothendieck category A is again a Grothendieck category.

Proof. Colimits and limits are taken dimensionwise, filtered colimits are obviously exact. Following notation of Hovey [13], denote by D^{n}X, n ∈ ℤ and X ∈ A, the complex which is X in degree n and n − 1 and 0 elsewhere, with interesting differential being the identity map. If U is a generator of A, then {D^{n}U}_{n∈ℤ} are generators of Ch(A). To see that these generate Ch(A), use the adjunction relation Hom_{Ch(A)}(D^{n}U,X) ≅ Hom_{A} (U,X_{n}). □

Remark 3.5. This remark is to warn the reader that one should not confuse generators in abelian and triangulated categories. Precisely, we shall also work with the derived category D(A) of unbounded complexes of a Grothendieck category A. Then generators for D(A) cannot be generators for Ch(A) and vice versa in general. Indeed, the generators {D^{n}U}_{n∈ℤ} of Ch(A) are contractible complexes, and hence zero in D(A).

On the other hand, suppose U is a generator for A. Denote by S^{n}U, n ∈ ℤ, the complex which is U in degree n and 0 elsewhere. Then {S^{n}U}_{n∈ℤ} is a family of generators for the derived category D(A) in the sense that for every non-zero object X ∈ D(A) there is a nonzero morphism in D(A) from some S^{n}U to X. But these cannot generate Ch(A) as the following example shows.

Suppose K is a field, and

\xymatrix{··· \ar[r] & 0 \ar[r] \ar[d] & K \ar[r] \ar[d]_{α} & 0 \ar[d] \ar[r] & 0 \ar[d] \ar[r] & ··· \\ ··· \ar[r] & 0 \ar[r] \ar[d] & K ⊕ K \ar[r]^{d_{0}} \ar[d]_{f} & K \ar[r] \ar[d] & 0 \ar[d] \ar[r] & ··· \\ ··· \ar[r] & 0 \ar[r] & K \ar[r] & 0 \ar[r] & 0 \ar[r] & ···}

is a commutative diagram in Ch(ModK), where d_{0}(x,y) = f(x,y) = y. We suppose the middle complex is concentrated in degrees 0 and −1. Clearly α(1) ∈ Kerd_{0} implies α(1) = (x′,0) for some x′∈ K. But fα(1) = f(x′,0) = 0, so fα = 0. Thus there is no non-zero map from S^{0}K to the middle complex such that the composite fα ≠ 0. Since there is no non-zero morphism from S^{n}K to the middle complex such that fα ≠ 0 for any n ≠ 0, we see that {S^{n}K}_{n∈ℤ} are not generators for Ch(ModK).

Definition 3.6. Recall that an object A ∈ C is finitely generated if whenever there are subobjects A_{i} ⊆ A for i ∈ I satisfying A = ∑_{i∈I} A_{i}, then there is a finite subset J ⊂ I such that A = ∑_{i∈J} A_{i}. The category of finitely generated subobjects of C is denoted by fgC. The category is locally finitely generated provided that every object X ∈ C is a directed sum X = ∑_{i∈I} X_{i} of finitely generated subobjects X_{i}, or equivalently, C possesses a family of finitely generated generators.

Definition 3.7. A finitely generated object B ∈ C is finitely presented provided that every epimorphism η : A → B with A finitely generated has a finitely generated kernel Kerη. The subcategory of finitely presented objects of C is denoted by fpC. The corresponding categories of finitely presented left and right R-modules over the ring R are denoted by Rmod= fp(RMod) and modR = fp(ModR), respectively. Note that the subcategory fpC of C is closed under extensions. Moreover, if

0 → A → B → C → 0

is a short exact sequence in C with B finitely presented, then C is finitely presented if and only if A is finitely generated. The category is locally finitely presented provided that it is locally finitely generated and every object X ∈ C is a direct limit X = \underrightarrow{lim}_{i∈I} X_{i} of finitely presented objects X_{i}, or equivalently, C possesses a family of finitely presented generators.

Definition 3.8. A finitely presented object C of a locally finitely presented Grothendieck category C is coherent if every finitely generated subobject B of C is finitely presented. Equivalently, every epimorphism h: C → A with A finitely presented has a finitely presented kernel. Evidently, a finitely generated subobject of coherent object is also coherent. The subcategory of coherent objects of C is denoted by cohC. The category C is locally coherent provided that it is locally finitely presented and every object X ∈ C is a direct limit X = \underrightarrow{lim}_{i∈I} X_{i} of coherent objects X_{i}, or equivalently, C possesses a family of coherent generators.

The subcategories consisting of finitely generated, finitely presented and coherent objects are ordered by inclusion as follows:

C ⊇ fgC ⊇ fpC ⊇ cohC .

Examples 3.9. (1) The categories ModR and (B,Ab) (see Examples 3.2-3.3) are locally finitely presented. Representable functors {h^{B}}_{B∈B} are projective finitely presented generators of (B,Ab) (see [12] for details). The category Ch(A) with A a Grothendieck category (see Proposition 3.4) is locally finitely presented whenever so is A. If {U_{i}}_{i∈I} are finitely presented generators of A, then {D^{n}U_{i}}_{n∈ℤ,i∈I} are finitely presented generators of Ch(A).

(2) ModR is locally coherent if and only if the ring R is right coherent. (B,Ab) is locally coherent whenever B is closed under cokernels (see [12] for details). In turn, Ch(A) with A a Grothendieck category is locally coherent whenever so is A.

4. Grothendieck categories of enriched functors

In this section we prove that the category of enriched functors [C ,V ] is a Grothendieck category whenever V is a closed symmetric monoidal Grothendieck category, giving us new Grothendieck categories in practice. Also, localization theory of Grothendieck categories becomes available for [C ,V ]. Moreover, [C ,V ] is closed symmetric monoidal whenever C is a symmetric monoidal V -category.

Here are some examples of closed symmetric monoidal Grothendieck categories.

Examples 4.1. (1) Given any commutative ring R, the triple (ModR,⊗_{R},R) is a closed symmetric monoidal Grothendieck category.

(2) More generally, let X be a quasi-compact quasi-separated scheme. Consider the category Qcoh(O_{X}) of quasi-coherent O_{X}-modules. By [15, 3.1] Qcoh(O_{X}) is a locally finitely presented Grothendieck category, where quasi-coherent O_{X}-modules of finite type form a family of finitely presented generators. The tensor product on O_{X}-modules preserves quasi-coherence, and induces a closed symmetric monoidal structure on Qcoh(O_{X}).

(3) Let R be any commutative ring. Let C′= {C′_{n},∂′_{n}} and C′′= {C′′_{n},∂′′_{n}} be two chain complexes of R-modules. Their tensor product C′⊗_{R}C′′= {(C′⊗_{R}C′′)_{n},∂_{n}} is the chain complex defined by

(C′⊗_{R}C′′)_{n} = ⨁_{i+j=n} (C′_{i}⊗_{R}C′′_{j}),

and

∂_{n}(t′_{i}⊗ s′′_{j}) = ∂′_{i}(t′_{i}) ⊗ s′′_{j}+ (−1)^{i}t′_{i}⊗ ∂′′_{j}(s′′_{j}), for all t′_{i}∈ C′_{i}, s′′_{j}∈ C′′_{j}, (i + j = n),

where C′_{i}⊗_{R}C′′_{j}denotes the tensor product of R-modules C′_{i}and C′′_{j}. Then the triple (Ch(ModR),⊗_{R},R) is a closed symmetric monoidal Grothendieck category. Here R is regarded as a complex concentrated in zeroth degree.

(4) (ModkG,⊗_{k},k) is closed symmetric monoidal Grothendieck category, where k is a field and G is a finite group.

The main result of this section is as follows.

Theorem 4.2. Let V be a closed symmetric monoidal Grothendieck category with a set of generators {g_{i}}_{I}. If C is a small V -category, then the category of enriched functors [C ,V ] is a Grothendieck V -category with the set of generators {V (c,−)⊘g_{i} | c ∈ ObC ,i ∈ I}. Moreover, if C is a small symmetric monoidal V -category, then [C ,V ] is closed symmetric monoidal with monoidal product and internal Hom-object computed by formulas of Day (7) and (8).

Proof. If C is a small V -category, then [C ,V ] is a V -category (see p. 6). The internal Homobject is given by (6). Let show that [C ,V ] is a preadditive category. Given V -functors X,Y ∈ [C ,V ], we have that

Hom_{[C ,V ]}(X,Y) = Hom_{V}(e,∫_{c∈ObC} V (X(c),Y (c))

is an abelian group, because V is preadditive. We can also describe explicitly the abelian group structure as follows. The morphisms of [C ,V ] are, by definition, the V -functors from C to V . Using Corollary 2.5, for any V -natural transformations

α,α′: X → Y

its sum α + α′is determined by the arrows

α_{c}+ α′_{c}: e → V (X(c),Y (c)).

Recall that Hom_{V}(X(c),Y (c)) = Hom_{V}(e,V (X(c),Y (c))) and α_{c}+α′_{c}is addition of α_{c} and α′_{c} in the abelian group Hom_{V}(X(c),Y (c)).

To show that the addition is bilinear, let

β ∈ Hom_{V}(e,∫_{c}V (Y (c),Z(c)) =∫_{c}Hom_{V}(e,V (Y (c),Z(c))
= ∫_{c}Hom_{V}(Y (c),Z(c)).

Using Corollary 2.5, we set

(βα)_{c} : = β_{c}∘ α_{c}.

Using the fact that V is preadditive, we have

(β(α + α′))_{c} = β_{c}(α_{c}+ α′_{c})
= β_{c}∘ α_{c}+ β_{c}∘ α′_{c}
= (βα + βα′)_{c}.

Similarly, (α + α′)γ = αγ + α′γ. We see that [C ,V ] is preadditive.

Since V is a bicomplete closed symmetric monoidal category and C is a small V -category, then by Lemma 2.9 the category [C ,V ] is bicomplete. Moreover, limits and colimits are formed objectwise. In particular, [C ,V ] has finite products. It follows from [18, VIII.2.2] that [C ,V ] is an additive category. Furthermore, [C ,V ] has kernels and cokernels which are defined objectwise.

Given a morphism α in [C ,V ], the canonical map

ᾱ : Coker(kerα) → Ker(cokerα)

is an isomorphism objectwise. Corollary 2.6 implies ᾱ is an isomorphism. It follows that [C ,V ] is an abelian category.

Next, direct limits exist in [C ,V ] and are defined objectwise. They are exact in [C ,V ], because so are direct limits in V (by assumption, V is a Grothendieck category). So, [C ,V ] is an Ab5-category.

It remains to find generators for [C ,V ]. By [5, 2.4] [C ,V] is a closed V -module, and hence there is an action

⊘: [C ,V ]⊗ V → [C ,V].

Now for any non-zero functor X ∈ [C ,V ]we have natural isomorphisms

Hom_{[C ,V ]}(V (c,−) ⊘ g_{i},X)) ≅ Hom_{V}(g_{i},V_{[C ,V ]}(V (c,−),X))
≅ Hom_{V}(g_{i},X(c)).

Let α : X → Y be a non-zero map in [C ,V ]. We want to show that there are i ∈ I, a map β : V (c,−) ⊘ g_{i} → X such that αβ ≠ 0.

Since α is non-zero, then α_{c} : X(c) → Y (c) is a non-zero map in V for some c ∈ ObC. By assumption {g_{i}}_{I} are generators of V , and so there is a map β̄: g_{i} → X(c) such that α_{c}β̄ ≠ 0. By the above isomorphism we can find a unique map β : V (c,−)⊘g_{i} → X corresponding to β̄. Now α_{c}β̄ ≠ 0 implies αβ ≠ 0 as required.

If C is a small symmetric monoidal V -category, then [C ,V ] is closed symmetric monoidal by Day’s Theorem 2.11. Monoidal product and internal Hom-object are computed by formulas of Day (7) and (8). □

Below we give a couple of examples illustrating the preceding theorem.

Example 4.3. Let R be a commutative ring with unit. Consider the closed symmetric monoidal category V = (ModR,⊗_{R},R). Consider a V -category C defined as follows. Its objects are integers ObC = ℤ. Given two integers m,n ∈ ObC, we define a Hom-object as

V_{C}(m,n) = \{ \begin{matrix}0 & if m ≠ n; \\ R & if m = n. \end{matrix}

Clearly, C is a V -category and the category [C ,V ] of V -functors from C to V is the product category ∏ _{ℤ} ModR. By definition, Ob(∏ _{ℤ} ModR) are tuples (M_{i})_{i∈ℤ} and morphisms are tuples of R-homomorphisms (f_{i} : M_{i} → N_{i})_{i∈ℤ}.

Let GrR be the category of ℤ-graded R-modules and graded homomorphisms. It is easy to see that the functor

∏_{ℤ} ModR → GrR, (M_{i})_{i∈ℤ} ↦ ⊕_{ℤ}M_{i},

is an isomorphism of categories. It is well known that GrR is a closed symmetric monoidal Grothendieck category with tensor product

(M ⊗_{R}N)_{k}:= ⨁_{i+j=k} M_{i}⊗_{R}N_{j} for all M,N ∈ GrR.  (9)

We want to show that this tensor product is recovered from Day’s theorem for [C ,V ]. Indeed, we define a symmetric monoidal product on C as follows. For every m,n ∈ ObC

m ⊠ n:= m + n.

Given a ∈ V_{C}(m,m) = R and b ∈ V_{C}(n,n) = R, we set

a ⊠ b = a · b ∈ V_{C}(m + n,m + n) = R.

Clearly, m ⊠ n = n ⊠ m. Since R is commutative, it follows that ⊠ defines a strictly symmetric monoidal tensor product on C.

Now for every M,N ∈ [C ,V ] ≅ GrR, Day’s theorem implies

M ⊙ N = ∫ ^{(m,n)∈ℤ⊠ℤ}M(m) ⊗_{R}N(n) ⊗_{R}V_{C}(m ⊠ n,−)
= ∫ ^{(m,n)∈ℤ⊠ℤ}M_{m}⊗_{R}N_{n}⊗_{R}V_{C}(m + n,−).

Thus,

(M ⊙ N)_{k} = ⨁_{m+n=k} M_{m}⊗_{R}N_{n}

and

\underline{Hom}(M,N)(n) =∫_{m∈ObC=ℤ} Hom_{R}(M(m),N(m ⊠ n))
= ∫_{m}Hom_{R}(M_{m},N_{m+n}) = Hom_{GrR}(M,N(n)).

So tensor product (9) as well as internal Hom-functor for graded modules are recovered from Day’s theorem.

By Theorem 4.2 [C ,V ] is Grothendieck with {R(n)}_{n∈ℤ} generators, where

R(n)_{m}= \{ \begin{matrix}0, & if m ≠ n; \\ R, & if m = n. \end{matrix}

It follows that GrR is Grothendieck. Moreover, [C ,V ] ≅ GrR is closed symmetric monoidal.

Example 4.4. Let R be a commutative ring. Recall that an R-linear DG-category C is just a category enriched over V = Ch(ModR). Right DG-modules over the DG-category C are just contravariant V -functors from C to V . In turn, DG-morphisms of DG-modules are nothing but V -natural transformations. The category of right DG-modules and DG-morphisms is denoted by ModC. Using our notation, one has, by definition, ModC = [C^{op},V ].

Theorem 4.2 and Proposition 3.4 imply ModC is a Grothendieck category with the family of generators

D^{n}(V_{C}(−,c)) := V_{C}(−,c) ⊘ D^{n}R, c ∈ ObC ,n ∈ ℤ.

Here D^{n}R stands for the complex which is R in degrees n and n − 1 and zero elsewhere, with interesting differential being the identity.

Example 4.5. Any preadditive category B is nothing but a category enriched over abelian groups V = Ab. V -functors from B to V are the same as additive functors. Theorem 4.2 says that the category of additive functors (B,Ab) is Grothendieck with representable functors {h^{B}= V_{B} (B,−) ⊘ ℤ}_{B∈B} being a family of generators. Thus the fact that (B,Ab) is Grothendieck (see Example 3.3) follows from Theorem 4.2.

5. Localization with respect to enriched subcategories

We say that a full subcategory S of an abelian category C is a Serre subcategory if for any short exact sequence

0 → X → Y → Z → 0

in C an object Y ∈ S if and only if X, Z ∈ S. A Serre subcategory S of a Grothendieck category C is localizing if it is closed under taking direct limits. Equivalently, the inclusion functor i: S → C admits the right adjoint t = t_{S} : C → S which takes every object X ∈ C to the maximal subobject t(X) of X belonging to S. The functor t we call the torsion functor. An object C of C is said to be S -torsionfree if t(C) = 0. Given a localizing subcategory S of C the quotient category C /S consists of C ∈ C such that t(C) = t^{1}(C) = 0, where t^{1} stands for the first derived functor associated with t. The objects from C /S we call S -closed objects. Given C ∈ C there exists a canonical exact sequence

0 → A′→ C −→^{λ_{C}} C_{S} → A′′→ 0

with A′= t(C), A′′∈ S, and where C_{S} ∈ C /S is the maximal essential extension of C̃ = C/t(C) such that C_{S} /C̃∈ S. The object C_{S} is uniquely defined up to a canonical isomorphism and is called the S -envelope of C. Moreover, the inclusion functor ι : C /S → C has the left adjoint localizing functor (−)_{S} : C → C /S, which is also exact. It takes each C ∈ C to C_{S} ∈ C /S. Then,

Hom_{C}(X,Y) ≅ Hom_{C /S} (X_{S} ,Y)

for all X ∈ C and Y ∈ C /S.

If C and D are Grothendieck categories, q: C → D is an exact functor, and a functor s: D → C is fully faithful and right adjoint to q, then S := Kerq is a localizing subcategory and there

exists an equivalence C /S ≅^{H} D such that H ∘ (−)_{S} = q. We shall refer to the pair (q,s) as the localization pair.

Example 5.1. Let A be a small preadditive category. Consider the category (A ,Ab) of additive functors from A to Ab. Let p ∈ ObA, then we have

Hom_{(A ,Ab)}((p,−),(p,−)) = End_{A} p.

Let

S_{p} = {F ∈ (A ,Ab) | F (p) = 0}.

Then S_{p} is a localizing subcategory of (A ,Ab). By [6] and [9] there is an equivalence of categories

(A ,Ab)/S_{p} ≅ Mod(End_{A} p)^{op}.

This result has some applications in ring and module theory (see [7, 8]).

More generally, given a collection of objects P in A, we can consider the localizing subcategory

S_{P} = {F ∈ (A ,Ab) | F (p) = 0 for all p ∈ P }

of (A ,Ab) and then

(A ,Ab)/S_{P} ≅ (P,Ab),

where P on the right hand side is regarded as a full subcategory in A (see [6] for details).

Our next goal is to obtain an enriched analog of this result.

Suppose V is a closed symmetric monoidal Grothendieck category. Let C be a V -category. By Theorem 4.2 [C ,V ] is a Grothendieck category. Suppose P is a collection of objects in C. We shall also regard P as a natural V -subcategory. Then

S_{P} = {F ∈ [C ,V ] | F (p) = 0 for all p ∈ P }

is localizing in [C ,V ].

We shall prove below that there is a natural equivalence of Grothendieck categories

[C ,V ]/S_{P} ≅ [P,V ].

Thus the same result of [6, 9] for the category of additive functors (A ,Ab) with A preaddtive (see Example 5.1 above) is recovered from the case when V = Ab. But first we prove the following

Proposition 5.2. Suppose V is a closed symmetric monoidal Grothendieck category. Let C be a V -category and let P consist of a collection of objects of C. Then the inclusion map i: P → C is a V -functor. It induces two adjoint functors

i_{∗} : [P,V ] ⇄ [C ,V ]: i^{∗}

where i_{∗} is the enriched left Kan extension and i^{∗} is just restriction to P.

Proof. Although this fact is a consequence of [2, 6.7.7], we give a proof here for the convenience of the reader.

If F ∈ [P,V ] then by Corollary 2.10 we have

F ≅ ∫ ^{ObP} V (p,−) ⊘ F (p).

By definition of the left Kan extension, we have

i_{∗}F ≅ ∫ ^{ObP} V (i(p),−) ⊘ F (p).

We want to show that

Hom_{[C ,V ]}(i_{∗}F,G) ≅ Hom_{[P,V ]}(F,i^{∗}G).

Using (6), one has isomorphisms of V -objects

V_{F} (F,i^{∗}G) = V_{F} (F,G ∘ i) = ∫_{ObP} V (F (p),G(i(p))).  (10)

On the other hand,

V_{F} (i_{∗}F,G)= V_{F} ( ∫ ^{ObP} V (i(p),−) ⊘ F (p),G)
= ∫_{ObP} V_{F} (F (p),V_{[C ,V ]}(V ((i(p),−),G)))
= ∫_{ObP} V (F (p),G(i(p))).  (11)

We have used here the fact that the functor V_{F} (−,G) takes V -coends to V -ends [2, 6.6.11] as well as the fact that [C ,V ] is a closed V -module. Now (10) and (11) imply i_{∗} and i^{∗} are adjoint functors. □

Theorem 5.3. Let S_{P} = {G ∈ [C ,V ]| G(p) = 0 for all p ∈ P}. Then S_{P} is a localizing subcategory of [C ,V ]and [P,V ] is equivalent to the quotient category [C ,V ]/S_{P} .

Proof. Obviously, S_{P} is localizing. Let

κ : [P,V ] → [C ,V ]/S_{P}

be the composition of the left Kan extension functor i_{∗} : [P,V ] → [C ,V ]and the localization functor (−)_{S_{P}} : [C ,V ]→ [C ,V ]/S_{P} . We want to prove that κ is an equivalence of categories.

First observe that

i^{∗}i_{∗}F ≅ F

for all F ∈ [P,V ]. Second, given G ∈ [C ,V ]the adjunction map β : i_{∗}i^{∗}G → G is such that Kerβ,Cokerβ ∈ S_{P} . Indeed, applying the exact functor i^{∗} to the exact sequence

[figure: Kerβ; i_{∗}i^{∗}G; β; G; Cokerβ,]

we get an exact sequence in [P,V ]

[figure: i^{∗}(Kerβ); i^{∗}i_{∗}i^{∗}G; i^{∗}(β); i^{∗}G; i^{∗}(Cokerβ).]

Since the composite map

i^{∗}G → i^{∗}i_{∗}i^{∗}G −⟶^{i^{∗}(β)} i^{∗}G

is the identity map and the left arrow is an isomorphism, then so is the right arrow. Thus

i^{∗}(Kerβ) = i^{∗}(Cokerβ) = 0,

and hence Kerβ,Cokerβ ∈ S_{P} . It also follows that

(i_{∗}i^{∗}G)_{S_{P}} ≅ G_{S_{P}} .  (12)

We have for all F,F′∈ [P,V ]

Hom_{[C ,V ]/S_{P}} (κ(F),κ(F′))≅ Hom_{[C ,V ]/S_{P}} ((i_{∗}(F)_{S_{P}} ,i_{∗}(F′)_{S_{P}} )

≅ Hom_{[C ,V ]}(i_{∗}F,(i_{∗}F′)_{S_{P}} )

≅ Hom_{[P,V ]}(F,i^{∗}((i_{∗}F′)_{S_{P}} ))

≅ Hom_{[P,V ]}(F,i^{∗}i_{∗}F′)

≅ Hom_{[P,V ]}(F,F′).

We use here an isomorphism i^{∗}G ≅ i^{∗}(G_{S_{P}} ) for any G ∈ [C ,V ]. The isomorphism is obtained by applying the exact functor i^{∗} to the exact sequence

[figure: S; G; λ_{G}; G_{S_{P}}; S′,; S,S′∈ S_{P} ,]

where λ_{G} is the S_{P} -envelope of G.

We see that κ is fully faithful. Now let G ∈ [C ,V ]/S_{P} be any S_{P} -closed object, then isomorphism (12) implies

κ(i^{∗}G) = (i_{∗}i^{∗}G)_{S_{P}} ≅ G_{S_{P}} ≅ G.

If we set F := i^{∗}G, then κ(F) ≅ G. This shows that κ is an equivalence of categories. □

Corollary 5.4 ([6]). Let C be a Grothendieck category with finitely generated projective generators B = {p_{i}}_{i∈I}. Let P = {p_{j}}_{j∈J} be a subfamily in B, where J ⊂ I. Then

S_{P} = {x ∈ C |Hom_{C}(p_{j},x) = 0 for all p_{j} ∈ P}

is localizing and C /S_{P} is a Grothendieck category with {(p_{j})_{S_{P}} } a family of finitely generated projective generators.

Proof. By Mitchell’s Theorem C is equivalent to the category (B^{op},Ab) by means of the functor T sending x ∈ C to (−,x). Now the latter category is the same as the category of enriched V - functors from B^{op} to V where V = Ab. We use as well the fact that a category is preadditive if and only if it is enriched in Ab.

It follows that T induces an equivalence of categories C /S_{P} and (B^{op},Ab)/S̃_{P} , where

S̃_{P} = {F ∈ (B^{op},Ab)|F (p_{j}) = 0 for all p_{j} ∈ P}.

Theorem 5.3 implies the latter quotient category is equivalent to (P^{op},Ab). The proof of Theorem 5.3 shows that the functor

(P^{op},Ab) → C /S_{P}

which sends F ∈ (P^{op},Ab) to (T^{−1} (i_{∗}F))_{S_{P}} is an equivalence of categories. It follows that {(p_{j})_{S_{P}} }_{j∈J} is a family of finitely generated projective generators of C /S_{P} . □

Example 5.5. Let C and V be as in Example 4.3 and let n ∈ ObC = ℤ. We set

S_{n} = {F ∈ [C ,V ] ≅ GrR | F (n) = 0}.

Then Theorem 5.3 implies

[C ,V ]/S_{n} ≅ GrR/S_{n} ≅ [P,ModR],

where P has one object n ∈ ℤ and V_{P} (n,n) = R. But [P,ModR] = ModR, hence GrR/S_{n} ≅ ModR.

Lemma 5.6. Let R be a ring object of a closed symmetric monoidal Grothendieck category V . Then the category of left R-modules R Mod can naturally be identified with the Grothendieck category [C ,V ], where ObC = {∗} and V_{C}(∗,∗) = R.

Proof. Given a left R-module M, define a V -functor F : C → V as follows. We set F (∗) = M and the structure map

F : V (∗,∗) → V (F (∗),F (∗))

to be the map adjoint to the structure map R ⊗ M → M. This rule clearly yields the desired identification. □

Example 5.7. To illustrate the previous lemma, let R be a commutative unital ring and let A = A_{0}⊕A_{1}⊕···be a graded R-algebra. Then A is a ring object in [C ,V ] = GrR with respect to tensor product (9). We can regard A as a one object V -category, where V = GrR. In our notation A = V_{GrR}(∗,∗). Lemma 5.6 implies [C ,V ] ≅ AMod is a Grothendieck category. Moreover, {A(n) = A ⊘ R(n)}_{n∈ℤ} are generators by Theorem 4.2 (see as well Example 4.3).

Corollary 5.8. Let C be a small V -category, and let c be any object of C. Then there is a natural equivalence of Grothendieck categories

RMod ≅ [C ,V ]/S_{c},

where S_{c} = {G ∈ [C ,V ]| G(c) = 0} and R = V (c,c).

Proof. This follows from Theorem 5.3 and Lemma 5.6. □

6. Comparing ChC_{R} and [modR,Ch(ModR)]

In the remaining sections we prove that the category of chain complexes of generalized modules ChC_{R} over a commutative ring R can be identified with the category of enriched functors [modR,Ch(ModR)], where the category of finitely presented modules modR is regarded as a full subcategory of Ch(ModR) of complexes concentrated in zeroth degree and this single entry is finitely presented. As an application of this, we show that ChC_{R} is a closed symmetric monoidal category. We shall also establish that ChC_{R} is a closed symmetric monoidal model category with nice finiteness conditions. As another application, the derived category D(C_{R}) will be shown to be closed symmetric monoidal compactly generated triangulated category with duality on compact objects. However, compact objects are not strongly dualisable as it will be shown below. Thus D(C_{R}) is an example of a category which satisfies all the axioms of a unital algebraic stable homotopy theory in the sense of Hovey–Palmieri–Strickland [14] except the property that compact objects are strongly dualisable. This kind of categories is new for the authors. We refer to these as unital algebraic almost stable homotopy theories, a basic example of which is the category D(C_{R}).

Let Ch(ModR) be the category of chain complexes of modules over a commutative ring R. It is a closed symmetric monoidal cofibrantly generated and weakly finitely generated model category by [13, sections II.3 and IV.2]. Weak equivalences are the quasi-isomorphisms and fibrations are the surjective chain maps. The tensor product and internal Hom-object are defined as above. The monoidal unit is the chain complex with the ring R concentrated in zeroth degree and other degrees are zero.

The category of finitely presented modules modR is a small symmetric monoidal category naturally enriched over ModR. Moreover, modR can also be enriched over ChModR if we regard a module M ∈ modR as the chain complex with M concentrated in zeroth degree and other degrees are zero. Given, M,N ∈ modR the internal Hom-object is the chain complex

···→ 0 → Hom_{R}(M,N) → 0 → ···

where Hom_{R}(M,N) is concentrated in zeroth degree. Observe that this complex equals \underline{Hom}(M,N), where M,N are regarded as complexes defined above.

If there is no likelihood of confusion, we shall also write the internal Hom-chain complex as V (A_{•},B_{•}), where A_{•},B_{•} are two chain complexes. In other words,

V (A_{•},B_{•}) := \underline{Hom}(A_{•},B_{•}).

Following notation of section 2 we denoted by [modR,Ch(ModR)] the category of enriched functors from the small symmetric monoidal category modR to Ch(ModR) (see Definition 2.2).

Following Herzog [12], we define the category of generalized modules C_{R} as

C_{R}:= (modR,Ab),

whose objects are the additive functors F : modR → Ab from the category of right finitely presented R-modules modR to the category of abelian groups Ab. Its morphisms are the natural transformations of functors. Similarly, the category _{R}C consists of the additive functors from the category of left finitely presented R-modules to Ab.

Theorem 6.1. Suppose R is a commutative ring. Then the category of generalized R-modules C_{R} can naturally be identified with the category [modR,ModR].

Proof. We briefly recall the proof from [3] for the convenience of the reader.

I. We first associate to any object in [modR,ModR] an object in C_{R}.

Let F ∈ Ob[modR,ModR]. By Definition 2.2 F takes Ob(modR) to Ob(ModR) and for all M,M′∈ modR there is a R-module homomorphism

F_{MM′}: Hom_{R}(M,M′) → Hom_{R}(F (M),F (M′)).  (13)

Let f : M → M′ be a homomorphism in modR, then we set F (f) to be the image of f in Hom_{R}(M,M′) under (13). Observe that an R-module structure on F (M) is given by

x · r = F_{MM}(r)(x)

for all x ∈ F (M) and r ∈ R. Here F_{MM}(r) stands for the image of the right multiplication endomorphism r: M → M.

II. Next, we want to show that the morphisms of [modR,ModR] can naturally be regarded as morphisms of C_{R}.

We have to verify that V -natural transformations in [modR,ModR] are morphisms in C_{R}. Given F,G ∈ [modR,ModR], the first step shows that F,G ∈ C_{R}. So given a V -natural transformation t in [modR,ModR], we want to prove that t yields a morphism in C_{R} in a natural way. For t we have structure homomorphisms in ModR (see Definition 2.3)

t_{M}: R → Hom_{R}(F (M),G(M)).

for all M,M′∈ modR. For any M ∈ ModR, set

τ_{M}:= t_{M}(1) : F (M) → G(M).

Therefore t yields a natural transformation in C_{R}

τ : F → G.

We see that morphisms in [modR,ModR] can naturally be regarded as morphisms of C_{R}. So F (M) ∈ ModR.

III. In this step we shall show that any object F of C_{R} can be regarded as an enriched functor in [modR,ModR].

For any M ∈ modR we have that F (M) ∈ Ab. Let us show that F (M) is an R-module. For all x ∈ F (M), we have to define x · r, where r ∈ R. The element r defines an R-module homomorphism r: M → M sending m ∈ M to m · r. We have a morphism F (r) : F (M) → F (M) and we set

x · r:= F (r)(x).

So F (M) ∈ ModR. Now we define an enriched functor associated with F. We therefore define a morphism in ModR

F_{MM′}: Hom_{R}(M,M′) → Hom_{R}(F (M),F (M′)), F_{MM′}(f) := F (f),

for all M, M′in modR.

Next we construct diagram (4). We have that

u_{M}: R → Hom_{R}(M,M)

is given by the right multiplication homomorphism. One has

F_{MM}(u_{M}(r)) = F (r)

for all r ∈ R, and hence the diagram

\xymatrix{R \ar[dr]_{u_{F (M)}} \ar[r]^{u_{M}} & Hom_{R}(M,M) \ar[d]^{F_{MM}} \\  & Hom_{R}(F (M),F (M))}

is commutative. The structure of an enriched functor for F, denoted by the same letter, is completed.

So every object in C_{R} can naturally be regarded as an enriched functor in [modR,ModR].

IV. In this step we shall show that morphisms in C_{R} (recall that these are natural transformations of additive functors) can naturally be regarded as V -natural transformations in [modR,ModR], i.e. as morphisms of [modR,ModR].

Let τ : F → G be any natural transformation in C_{R}. Then for each object M ∈ modR, there exists a homomorphism τ_{M}: F (M) → G(M) in Aband for each homomorphism f : M → M′in modR the diagram

\xymatrix{F (M) \ar[d]_{F (f)} \ar[r]^{τ_{M}} & G(M) \ar[d]^{G(f)} \\ F (M′) \ar[r]_{τ_{M}} & G(M′)}

is commutative.

By above we can regard F,G as enriched functors. We want to show that τ yields a V -natural transformation t between the V -functors F and G. For any M ∈ modR define a map

t_{M}: R → Hom_{R}(F (M),G(M))

as t_{M}(r) := τ_{M}∘F (r) = G(r)∘τ_{M}. By definition, t_{M}(1) = τ_{M}. Then the maps t_{M} yield V -natural transformations between the V -functors F and G. □

More generally, we have the following

Theorem 6.2. Suppose R is a commutative ring. Then the category of chain complexes of generalized R-modules Ch(C_{R}) can naturally be identified with the category [modR,Ch(ModR)].

Proof. Given M ∈ modR we want to describe chain morphisms of the form:

\xymatrix{⋯ \ar[r] & 0 \ar[r]^{0} \ar[d] & M \ar[r]^{0} \ar[d]_{α} & 0 \ar[r] \ar[d] & ⋯ \\ ⋯ \ar[r] & V (A_{•},B_{•})_{1} \ar[r]^{∂_{1}} & V (A_{•},B_{•})_{0} \ar[r]^{∂_{0}} & V (A_{•},B_{•})_{−1} \ar[r] & ⋯}

From the diagram we have ∂_{0}α = 0. This means ∂_{0}α(m) = 0, for all m ∈ M.

α(m) ∈ V (A_{•},B_{•})_{0} = ∏_{p∈ℤ} Hom_{R}(A_{p},B_{p})
α(m) = (α_{p}(m) : A_{p} → B_{p})_{p∈ℤ}
∂α_{p}(m) = 0 = ∂α_{p}(m) − α(m)_{p−1} ∂.

We get a commutative diagram

\xymatrix{⋯ \ar[r] & A_{p+1} \ar[r]^{∂_{p+1}} \ar[d]_{α_{p+1} (m)} & A_{p} \ar[r]^{∂_{p}} \ar[d]_{α_{p}(m)} & A_{p−1} \ar[r] \ar[d]^{α_{p−1} (m)} & ⋯ \\ ⋯ \ar[r] & B_{p+1} \ar[r]_{∂_{p+1}} & B_{p} \ar[r]_{∂_{p}} & B_{p−1} \ar[r] & ⋯}

This shows α(m) : A_{•} → B_{•} is a chain map.

I. Consider a V -functor F ∈ [modR,Ch(ModR)]. By definition, we have F (M) ∈ Ch(ModR) for any M in modR. We also have a map

F_{MM′}: V (M,M′) → V (F (M),F (M′)) ∈ Ch(ModR).

This is the same as a chain map

F_{MM′}: (⋯ → 0 → Hom_{R}(M,M′) → 0 → ⋯) −→ \underline{Hom}(F (M),F (M′)).

By above we have a chain map

F_{MM′}(f) : F (M) → F (M′)

for all f ∈ Hom_{R}(M,M′). It is directly verified that F_{MM}(id_{M}) = id_{F (M)} and F_{MM′′}(gf) = F_{M′M′′}(g)F_{MM′}(f) for any g ∈ Hom_{R}(M′,M′′) (see Definition 2.2).

If we observe that ChC_{R} is the same as the category (modR,Ch(Ab)) of additive functors from modR to Ch(Ab), it follows that the enriched functor F gives rise to an object in ChC_{R}. We denote this object by the same letter.

II. Now let α : F ⇒ G be a V -map in [modR,Ch(ModR)]. It consists of giving a chain map

α_{M}: (0 → R → 0) −→ V (F (M),G(M)) = \underline{Hom}(F (M),G(M)),

which is equivalent to giving a chain map α_{M}(r) : F (M) → G(M) for all r ∈ R, such that the following diagram commutes

[figure: (0 → Hom_{R}(M,M′) → 0); l{−1 \atop V (M,M′)}; r{−1 \atop V (M,M′)}; (0 → R → 0) ⊗ (0 → Hom_{R}(M,M′) → 0); α_{M}⊗G_{MM′}; (0 → Hom_{R}(M,M′) → 0) ⊗ (0 → R → 0); F_{MM′}⊗α_{M′}; V (F (M),G(M)) ⊗ V (G(M),G(M′)); c_{F (M)G(M)G(M′)}; V (F (M),F (M′)) ⊗ V (F (M′),G(M′)); c_{F (M)F (M′)G(M′)}; V (F (M),G(M′))]

The diagram implies the following:

(a) For all r ∈ R and f ∈ Hom_{R}(M,M′) the chain map c(α_{M}(r)⊗G_{MM′}(f)) : F (M) → G(M′) equals the chain map G_{MM′}(f) ∘ α_{M}(r).

(b) For all r ∈ R and f ∈ Hom_{R}(M,M′) the chain map c(F_{MM′}(f)⊗α_{M}(r)) : F (M) → G(M′)

equals α_{M′}(r) ∘ F_{MM′}(f). So we get a commutative diagram

\xymatrix{F (M) \ar[r]^{α_{M}(r)} \ar[d]_{F_{MM′}(f)} & G(M) \ar[d]^{G_{MM′}(f)} \\ F (M′) \ar[r]_{α_{M′}(r)} & G(M′)}

Thus α_{M}(r) : F (M) → G(M) gives rise to a morphism in (modR,Ch(Ab)) = ChC_{R}.

We shall associate to the enriched map α : F ⇒ G in [modR,Ch(ModR)] the map F → G in ChC_{R} given by the chain maps α_{M}(1) : F (M) → G(M), M ∈ modR. We denote the associated chain map by the same letter α.

III. By Theorem 6.1 the category Ch[modR,ModR] is identified with ChC_{R}. Let F_{•} ∈ Ch[modR,ModR], so we have a chain complex

\xymatrix{F_{•} = ⋯ \ar[r] & F_{n+1} \ar[r]^{∂_{n}+1} & F_{n} \ar[r]^{∂_{n}} & F_{n−1} \ar[r]^{∂_{n−1}} & ⋯}

with each F_{n} ∈ [modR,ModR] and each ∂_{n} being a V -natural transformation from modR to ModR. We have to associate a V -functor in [modR,Ch(ModR)] to F_{•}.

If M ∈ modR then

\xymatrix{F_{•}(M) = ⋯ \ar[r] & F (M)_{n+1} \ar[r]^{∂_{n+1,M}(1)} & F (M)_{n} \ar[r]^{∂_{n,M}(1)} & F (M)_{n−1} \ar[r]^{∂_{n−1,M}(1)} & ⋯}∈ Ch(ModR)

Also, for any map f : M → M′∈ modR we have that F_{•}(f) : F_{•}(M) → F_{•}(M′) is a chain map, because each square of the following diagram is commutative

\xymatrix{⋯ \ar[r] & F_{n+1} (M) \ar[r]^{∂_{n+1}} \ar[d]_{F_{MM′}(f)} & F_{n}(M) \ar[r]^{∂_{n}} \ar[d]_{F_{MM′}(f)} & F_{n−1} (M) \ar[r] \ar[d]^{F_{MM′}(f)} & ⋯ \\ ⋯ \ar[r] & F_{n+1} (M′) \ar[r]_{∂_{n+1}} & F_{n}(M′)_{n} \ar[r]_{∂_{n}} & F (M′)_{n−1} , \ar[r] & ⋯}

Thus F_{MM′}(f) is a chain map, and hence one gets a chain map

F_{MM′}: (⋯ → 0 → Hom_{R}(M,M′) → 0 → ⋯) −→ \underline{Hom}(F_{•}(M),F_{•}(M′)).

Since F_{MM}(id_{M}) = id_{F_{•}(M)} and F_{MM′′}(gf) = F_{M′M′′}(g)F_{MM′}(f) for all g ∈ Hom_{R}(M′,M′′), F_{•} yields a V -functor in [modR,Ch(ModR)] denoted by the same letter.

IV. Next let β : F_{•} → G_{•} be a chain map in Ch[modR,ModR], then we have the following commutative diagram

\xymatrix{⋯ \ar[r] & F_{n+1} \ar[r]^{∂_{n+1}} \ar[d]_{β_{n+1}} & F_{n} \ar[r]^{∂_{n}} \ar[d]_{β_{n}} & F_{n−1} \ar[r] \ar[d]_{β_{n−1}} & ⋯ \\ ⋯ \ar[r] & G_{n+1} \ar[r]_{∂_{n+1}} & G_{n} \ar[r]_{∂_{n}} & G_{n−1} \ar[r] & ⋯}

Note that each β_{n,M}: R → Hom_{R}(F_{n}(M),G_{n}(M)) is such that diagram (5) is commutative for it. So we are given maps β_{n,M}(r) : F_{n}(M) → G_{n}(M) for all r ∈ R.

One has a commutative diagram for all r ∈ R

\xymatrix{··· \ar[r] & F_{n+1} (M) \ar[r]^{∂_{n+1 ,M}(1)} \ar[d]_{β_{n+1,M}(r)} & F_{n}(M) \ar[r]^{∂_{n,M}(1)} \ar[d]_{β_{n,M}(r)} & F_{n−1} (M) \ar[r] \ar[d]_{βn,M(r)} & ··· \\ ··· \ar[r] & G_{n+1} (M) \ar[r]_{∂_{n+1 ,M}(1)} & G_{n}(M) \ar[r]_{∂_{n,M}(1)} & G_{n−1} (M)) \ar[r] & ···}

In particular, β_{M}(r) : F_{•}(M) → G_{•}(M) is a chain map.

Now we want to show that the diagram

\xymatrix{F_{•}(M) \ar[r]^{β_{M}(r)} \ar[d]_{F_{•}(f)} & G_{•}(M) \ar[d]^{G_{•}(f)} \\ F_{•}(M′) \ar[r]_{β_{M′}(r)} & G_{•}(M′)}  (14)

commutes. Commutativity of diagram (5) implies commutativity of the following diagram for all n ∈ ℤ:

\xymatrix{F_{n}(M) \ar[r]^{β_{n,M}(r)} \ar[d]_{F_{n}(f)} & G_{n}(M) \ar[d]^{G_{n}(f)} \\ F_{n}(M′) \ar[r]_{β_{n,M′}(r)} & G_{n}(M′)}

Commutativity of the latter together with the facts that β_{M}(r),G_{•}(f),F_{•}(f) are all chain maps is enough to check commutativity of (14) as shown in the diagram below

\xymatrix{ & F_{n+1} (M) \ar[rr]^{d_{n+1}} \ar[dd]_{β_{n+1,M}(r)} \ar[dl]_{F_{n+1} (f)} &  & F_{n}(M) \ar[dd]_{β_{n,M(r)}} \ar[dl]_{F_{n}(f)} \\ F_{n+1} (M′) \ar[rr]^{d_{n+1}} \ar[dd]_{β_{n+1,M′}(r)} &  & F_{n}(M′) \ar[dd]_{βn,M′(r)} \\  & G_{n+1} (M) \ar[rr]^{d_{n+1}} \ar[dl]_{G_{n+1} (f)} &  & G_{n}(M) \ar[dl]^{G_{n}(f)} \\ G_{n+1} (M′) \ar[rr]_{d_{n+1}} &  & G_{n}(M′)}

Thus we have constructed a V -natural map β : F_{•} → G_{•}.

It is now easily verified that associations given in steps I–IV yield the desired isomorphisms of categories [modR,Ch(ModR)] and ChC_{R}. □

We are now in a position to prove the main result of this section.

Theorem 6.3. Let R be a commutative ring, then Ch(C_{R}) is a left and right proper closed symmetric monoidal V -model category, where V = Ch(ModR), and the monoid axiom in the sense of Schwede–Shipley [23] is satisfied. The tensor product of two complexes F_{•},G_{•} ∈ Ch(C_{R}) is given by

F_{•}⊙ G_{•} = ∫ ^{(M,N)∈modR⊗modR}F_{•}(M) ⊗_{R}G_{•}(N) ⊗_{R}Hom_{R}(M ⊗_{R}N,−).  (15)

Here Hom_{R}(M ⊗_{R}N,−) is regarded as a complex concentrated in zeroth degree. The internal Hom-object is defined as

\underline{Hom}(F_{•},G_{•})(M) = ∫_{N∈modR} \underline{Hom}_{Ch(ModR)}(F_{•}(N),G_{•}(M ⊗_{R}N)).  (16)

Proof. By Theorem 6.2 we know that ChC_{R}can be identified with the category [modR,Ch(ModR)]. Note that modR is symmetric monoidal category enriched over Ch(ModR). Now formulas (15)- (16) as well as the fact that ChC_{R} is closed symmetric monoidal follow from Theorem 2.11.

It remains to show that ChC_{R} is a left and right proper closed symmetric monoidal V -model category, and the monoid axiom is satisfied. Since ChC_{R}can be identified with [modR,Ch(ModR)] by Theorem 6.2, it is enough to verify this for [modR,Ch(ModR)]. The category Ch(ModR) is a closed symmetric monoidal cofibrantly generated model category, where the weak equivalences are quasi-isomorphisms and the fibrations are surjective chain morphisms (see [13, 2.3.11] and [13, 4.2.13]). Moreover, Ch(ModR) is weakly finitely generated in the sense of [5, Section 3.1]. Also, Ch(ModR) satisfies the monoid axiom in the sense of [23]. It follows from [5, 4.2] that [modR,Ch(ModR)] is a weakly finitely generated model category, where fibrations and weak equivalences are defined objectwise. Furthermore, [modR,Ch(ModR)] is a monoidal V - enriched model category satisfying the monoid axiom by [5, 4.4], because modR is a symmetric monoidal category enriched over Ch(ModR). Finally, [5, 4.8] implies [modR,Ch(ModR)] is both left and right proper. □

Corollary 6.4. Let R be a ring object in Ch(C_{R}). Then the category of left R-modules is a cofibrantly generated model Grothendieck category. If R is a commutative ring object, then the category of R-modules is a cofibrantly generated, monoidal model category satisfying the monoid axiom, and the category of R-algebras is a cofibrantly generated model category.

Proof. This is a consequence of Lemma 5.6, Theorem 6.3 and [23, 4.1]. □

Example 6.5. A typical example of a ring object in Ch(C_{R}) is constructed as follows. Let C be a finitely presented R-module. We regard the representable functor (C,−) ∈ C_{R}as a complex in zeroth degree. Lemma 2.12 implies a natural isomorphism of representable functors

(C,−) ⊙ (C,−) ≅ (C ⊗_{R}C,−).

It follows that (C,−) is a ring object of Ch(C_{R}) if and only if C is an R-coalgebra. It is a commutative ring object if and only if C is a cocommutative R-coalgebra. By the previous corollary we have a Grothendieck model category of (C,−)-modules inside Ch(C_{R}).

7. Almost stable homotopy category structure for D(C_{R})

In this section we give an application of Theorem 6.3. Namely, we prove that the derived category D(C_{R}) of the Grothendieck category of generalized modules is a closed symmetric monoidal compactly generated triangulated category with duality on compact objects. However, compact objects are not strongly dualizable as it will be shown below. Thus D(C_{R}) is an example of a category which satisfies all the axioms of a unital algebraic stable homotopy theory in the sense of Hovey–Palmieri–Strickland [14] except the property that compact objects are strongly dualizable. We refer to these as unital algebraic almost stable homotopy categories. A basic example of this class of categories is D(C_{R}).

Definition 7.1. Let C be a triangulated category in which all set indexed direct sums exist. An object A of C is called compact if the canonial map

⨁_{i∈I} Hom(A,E_{i}) → Hom(A,⨁_{i∈I} E_{i})

is an isomorphism for any set of objects E_{i} in C and i ∈ I. A triangulated category C is compactly generated if there is a set S of compact objects with the following property, if for every object E ∈ C we have

Hom(A,E) = 0 for all A ∈ S implies E = 0.

Recall that the derived tensor product F_{•} ⊙^{L}G_{•} of two complexes in ChC_{R} is defined as F_{•}^{c} ⊙ G_{•}^{c}, where F_{•}^{c},G_{•}^{c} are cofibrant replacements of F_{•} and G_{•} respectively in the monoidal model category ChC_{R}. We start with the following Theorem 7.2. Let R be a commutative ring. Then the derived category D(C_{R}) of the Grothendieck category C_{R} is a compactly generated triangulated closed symmetric monoidal category, where formulas (15) and (16) yield the derived tensor product F_{•} ⊙^{L}G_{•} and derived internal Homobject R\underline{Hom}(F_{•},G_{•}). The compact objects of D(C_{R}) are the complexes isomorphic to bounded complexes of coherent functors in cohC_{R}.

Proof. By Theorem 6.3 Ch(C_{R}) is a left and right proper closed symmetric monoidal V -model category, where V = Ch(ModR). Hence the derived category D(C_{R}) is identified with the homotopy category Ho(Ch(C_{R})). But the latter category is a closed symmetric monoidal category by [13, 4.3.2] with derived tensor product and internal interal Hom-functors induced by (15) and (16) from Theorem 6.3.

Since C_{R}is a Grothendieck category with finitely generated projective generators {(M,−)}_{M∈modR}, then its derived category D(C_{R}) is a compactly generated triangulated category. The compact objects are those complexes which are isomorphic in D(C_{R}) to bounded complexes of representable functors. They are also called perfect complexes. Since every coherent functor C ∈ cohC_{R} has a resolution (see, e.g., [12, 2.1])

\xymatrix{0 \ar[r] & (L,−) \ar[r] & (K,−) \ar[r] & (M,−) \ar[r] & C \ar[r] & 0,}

where K,L,M ∈ modR, then every bounded complex of coherent functors is quasi-isomorphic to a bounded complex of representable functors. It follows that every bounded complex of coherent functors is quasi-isomorphic to a perfect complex. This finishes the proof. □

Remark 7.3. A monoidal unit in ChC_{R} and D(C_{R}) is (R,−) ≅ − ⊗_{R}R regarded as a complex concentrated in zeroth degree.

Definition 7.4. Let V be a closed symmetric monoidal additive category, with monoidal product x ⊗ y, unit e, and internal function objects V (x,y). An object x ∈ V is strongly dualizable if the natural map V (x,e) ⊗ y → V (x,y) is an isomorphism for all y. We shall also write x^{∨} to denote V (x,e).

It follows from [17, Theorem 7.1.6] that the functor

x ∈ V ↦ x^{∨} ∈ V

puts the full subcategory of strongly dualizable objects of V in duality.

We want to show below that D(C_{R}) has a duality on compact objects but these are not strongly dualizable in general. To this end we shall need a categorical duality

D: (cohC_{R})^{op} → coh_{R}C

of Auslander [1] and Gruson–Jensen [10] (see [12] as well) defined over any non-commutative ring R as follows. Given_{R}N ∈ Rmod, we have

(DC)(_{R}N) := Hom_{C_{R}}(C,− ⊗_{R}N).

If η : B → C is a morphism in cohC_{R}, then

D(η)_{N} : D(C)(_{R}N) → D(B)(_{R}N)

is defined to be Hom_{C_{R}}(η,− ⊗_{R}N). For M_{R}∈ modR and_{R}N ∈ Rmodwe have that

D(M_{R},−) ≅ M ⊗_{R}− and D(− ⊗_{R}N) ≅ (_{R}N,−).

We shall refer to this duality as the Auslander–Gruson–Jensen Duality.

Suppose now R is commutative. Then the category C_{R} is closed symmetric monoidal for the same reasons that ChC_{R} is. The monoidal product C ⊙ C′and internal Hom-object \underline{Hom}(C,C′) are computed by formulas, which are similar to (15)-(16). It can also be shown that the Auslander– Gruson–Jensen Duality D defined above is isomorphic to the internal Hom-functor

\underline{Hom}(−,(R,−)) ≅ \underline{Hom}(−,− ⊗_{R}R)

(we refer the reader to [3] for further details).

The following example shows that compact objects of D(C_{R}) are not strongly dualizable in general.

Example 7.5. There are objects C ∈ D(C_{R})^{c} and X ∈ D(C_{R}) such that the natural arrow

C^{∨} ⊙^{L}X → R\underline{Hom}(C,X)  (17)

is not an isomorphism, where

C^{∨} := R\underline{Hom}(C,− ⊗_{R}R) = R\underline{Hom}(C,(R,−)).

Let R = ℤ and M = N = ℤ_{2} ∈ modℤ. We have an exact sequence

\xymatrix{0 \ar[r] & ℤ \ar[r]^{2} & ℤ \ar[r] & ℤ_{2} \ar[r] & 0.}

We want to compute (− ⊗ M)^{∨} ⊙^{L}− ⊗ N = (− ⊗ ℤ_{2})^{∨} ⊙^{L}− ⊗ ℤ_{2}.

To compute (− ⊗ ℤ_{2})^{∨}, consider a projective resolution for − ⊗ ℤ_{2} in C_{ℤ}

\xymatrix{0 \ar[r] & (ℤ_{2},−) \ar[r] & (ℤ,−) \ar[r]^{(2,−)} & (ℤ,−) \ar[r] & − ⊗ ℤ_{2} \ar[r] & 0.}

Then (− ⊗ ℤ_{2})^{∨} is the value of \underline{Hom}(−,(ℤ,−)) at the projective resolution of − ⊗ ℤ_{2}. Since \underline{Hom}(−,(ℤ,−)) puts the category of coherent objects cohC_{ℤ} in duality and takes representable functors (L,−) to − ⊗ L, then (− ⊗ ℤ_{2})^{∨} is the complex

···→ 0 → − ⊗ ℤ → − ⊗ ℤ → − ⊗ ℤ_{2} → 0 → ···

This complex has only one non-zero homology group (ℤ_{2},−), which we place in zeroth degree as a complex. We see that (− ⊗ ℤ_{2})^{∨} ≅ (ℤ_{2},−).

We take a projective resolution for − ⊗ N in cohC_{ℤ} as above

\xymatrix{0 \ar[r] & (ℤ_{2},−) \ar[r] & (ℤ,−) \ar[r]^{(2,−)} & (ℤ,−) \ar[r] & − ⊗ ℤ_{2} \ar[r] & 0.}

Tensoring it with (M,−) = (ℤ_{2},−) we get (M,−) ⊙^{L}− ⊗ N which is the complex

···→ 0 → (ℤ_{2},−) ⊙ (ℤ_{2},−) → (ℤ_{2},−) ⊙ (ℤ,−) → (ℤ_{2},−) ⊙ (ℤ,−) → 0 → ···

By Lemma 2.12 it equals the complex

\xymatrix{··· \ar[r] & 0 \ar[r] & (ℤ_{2},−) \ar[r] & (ℤ_{2},−) \ar[r]^{0} & (ℤ_{2},−) \ar[r] & 0 \ar[r] & ···}

Note that evaluation of this complex at ℤ is zero.

Now compute R\underline{Hom}(− ⊗ M,− ⊗ N). It is the value of \underline{Hom}(−,− ⊗ N) at the projective resolution of − ⊗ M

\xymatrix{0 \ar[r] & (ℤ_{2},−) \ar[r] & (ℤ,−) \ar[r]^{(2,−)} & (ℤ,−) \ar[r] & − ⊗ M \ar[r] & 0.}

Applying \underline{Hom}(−,− ⊗ N) to the complex

\xymatrix{0 \ar[r] & (ℤ_{2},−) \ar[r] & (ℤ,−) \ar[r]^{(2,−)} & (ℤ,−) \ar[r] & 0}

we get a complex

0 → \underline{Hom}((ℤ,−),− ⊗ ℤ_{2}) → \underline{Hom}((ℤ,−),− ⊗ ℤ_{2}) → \underline{Hom}((ℤ_{2},−),− ⊗ ℤ_{2}) → 0

Using enriched Yoneda Lemma 2.8, it is equal to

\xymatrix{0 \ar[r] & − ⊗ ℤ_{2} \ar[r]^{0} & − ⊗ ℤ_{2} \ar[r]^{≅} & − ⊗ ℤ_{2} \ar[r] & 0.}

The value of this complex at ℤ has non-trivial homology. We conclude that (17) cannot be an isomorphism in general. We conclude that compact objects of D(C_{R}) are not strongly dualizable.

Lemma 7.6. The triangulated category D(C_{R})^{c} of compact objects of D(C_{R}) is triangle equivalent to the derived category D^{b}(cohC_{R}) of bounded complexes in cohC_{R}.

Proof. By the proof of Theorem 7.2 D(C_{R})^{c} is triangle equivalent to the full subcategory D̃(C_{R})^{c} of bounded complexes in cohC_{R}. Let

\xymatrix{ & X \ar[dl]_{∽} \ar[dr] \\ M &  & N}

be a morphism in D(C_{R}) with M,N ∈ D̃(C_{R})^{c}. Let P → M be a projective resolution of M. Then P is in D̃(C_{R})^{c} and P is isomorphic to X in D(C_{R}). But P is a bounded complex of projectives in C_{R}. Therefore there is a quasi-isomorphism P → X. Then we get a diagram in D̃(C_{R})^{c}

\xymatrix{ & P \ar[d]_{∽} \\  & X \ar[dl]_{∽} \ar[dr] \\ M &  & N}

By [16, 9.1] the natural functor

D^{b}(cohC_{R}) → D̃(C_{R})^{c}

is fully faithful. But objects are the same, and therefore these subcategories of D(C_{R}) coincide. □

Lemma 7.7. There is a triangle equivalence of triangulated categories D^{b}(cohC_{R}) and D^{b}((cohC_{R})^{op}) taking X ∈ D^{b}(cohC_{R}) to \underline{Hom}(X,− ⊗ R).

Proof. By [3] \underline{Hom}(−,− ⊗ R) : cohC_{R} → (cohC_{R})^{op} is an equivalence of abelian categories. Moreover, this functor is isomorphic to the Auslander–Gruson–Jensen duality (see [3, 4.6]). The fact that equivalent abelian categories have equivalent derived categories finishes the proof. □

Corollary 7.8. D^{b}(cohC_{R}) is triangle equivalent to (D^{b}(cohC_{R}))^{op}.

Proof. This follows from the previous lemma and the fact that D^{b}(A^{op}) is triangle equivalent to (D^{c}(A))^{op} for any abelian category A. □

We are now in a position to prove the following Theorem 7.9 (Auslander–Gruson–Jensen Duality for compact objects). Let D(C_{R})^{c} be the full triangulated subcategory of D(C_{R}) of compact objects. Then there is a duality

D: (D(C_{R})^{c})^{op} → D(C_{R})^{c}

that takes a compact object C_{•} to

DC_{•} := R\underline{Hom}(C_{•},− ⊗_{R}R).

Proof. By Lemma 7.6 D(C_{R})^{c} ≃ D^{b}(cohC_{R}). Let D^{b}(projC_{R}) be a full subcategory of D(C_{R})^{c} consisting of bounded complexes of representable coherent functors. The composition

\xymatrix{D^{b}(projC_{R}) \ar[r] & D(C_{R})^{c} \ar[r]^{≃} & D^{b}(cohC_{R})}

is an equivalence of triangulated categories.

Lemma 7.7 and Corollary 7.8 imply \underline{Hom}(−,− ⊗ R) is an equivalence of triangulated categories D(cohC_{R}) ≃ (D^{b}(cohC_{R}))^{op}.

Now the composite of equivalences

D^{b}(projC_{R}) → D(C_{R})^{c} → D^{b}(cohC_{R}) → (D^{b}(cohC_{R}))^{op} → (D(C_{R})^{c})^{op}

computes the desired equivalence R\underline{Hom}(−,− ⊗ R) of triangulated categories. □

Definition 7.10 (Hovey–Palmieri–Strickland [14]). A stable homotopy category is a category C with the following extra structure:

(1) A triangulation.

(2) A closed symmetric monoidal structure, compatible with the triangulation.

(3) A set G of strongly dualizable objects of C, such that the only localizing subcategory of C containing G is C itself. We also assume that C satisfies the following:

(4) Arbitrary coproducts of objects of C exist.

(5) Every cohomology functor on C is representable.

We shall say that such a category C is algebraic if the objects of G are compact. If, in addition, the unit object e is compact, we say that C is unital algebraic.

Definition 7.11. An almost stable homotopy category is a category C which satisfies axioms (1)-(2) and (4)-(5) for a stable homotopy category and the following axiom:

(3′) There is a full small subcategory G of C with duality D: G^{op} → G, such that the only

localizing subcategory of C containing G is C itself. Algebraic and unital algebraic almost stable homotopy categories are defined as in Definition 7.10.

Remark 7.12. Every stable homotopy category C with generating set of strongly dualizable objects G is an almost stable homotopy category. Indeed, we can assume without loss of generality that x^{∨} ∈ G for every x ∈ G, and then the full subcategory of C whose objects are those of G has a duality x ↦ x^{∨} and generates C. The following theorem shows that there are algebraic almost stable homotopy categories which are not stable homotopy categories.

Theorem 7.13. Let R be a commutative ring. Then D(C_{R}) is a unital algebraic almost stable homotopy category, which is not an algebraic stable homotopy category in the sense of Definition 7.10.

Proof. Let G be the full subcategory of compact objects of D(C_{R}). The fact that D(C_{R}) is a unital algebraic almost stable homotopy category follows from Theorems 7.2 and 7.9. We also use here the fact that every cohomology functor on a compactly generated triangulated category is representable by a theorem of Neeman [19, Theorem 3.1].

Suppose D(C_{R}) is generated by compact strongly dualizable objects G as required for an algebraic stable homotopy category. By [14, Theorem A.2.5] we may assume without loss of generality that G is a thick subcategory in the triangulated category of compact objects. If G generated D(C_{R}), then another theorem of Neeman [19, Theorem 2.1] would imply that G contains all compact objects. But Example 7.5 shows that compact objects of D(C_{R}) are not strongly dualizable in general. This contradiction finishes the proof. □

References

[1] M. Auslander, Coherent functors, Proceedings of the Conference on Categorical Algebra (La Jolla 1965) (eds S. Eilenberg, D. K. Harrison, S. MacLane, and H. Röhrl, Springer, New York, 1966), pp. 189-231.

[2] F. Borceux, Handbook of categorical algebra. 2, Cambridge University Press, Cambridge, 1994.

[3] J. Darbyshire, Monoidal structures for generalized modules, MPhil Project, Swansea University, 2013.

[4] B. Day, On closed categories of functors, In Reports of the Midwest Category Seminar, IV, pp. 1-38. Springer, Berlin, 1970.

[5] B. I. Dundas, O. Röndigs, P. A. Østvær, Enriched functors and stable homotopy theory, Doc. Math. 8 (2003), 409-488.

[6] G. A. Garkusha, Grothendieck categories, Algebra i Analiz 13(2) (2001), 1-68 (Russian). Engl. transl. in St. Petersburg Math. J. 13(2) (2002), 149-200.

[7] G. A. Garkusha, A. I. Generalov, Duality for categories of finitely presented modules, Algebra i Analiz 11(6) (1999), 139-152. (Russian). English transl. in St. Petersburg Math. J. 11(6) (2000), 1051-1061.

[8] G. A. Garkusha, FP-injective and weakly quasi-Frobenius rings, Zap. Nauchn. Sem. S.-Peterburg. Otd. Mat. Inst. Steklov. (POMI) 265 (1999), 110-129. (Russian). English transl. in J. Math. Sci. 112(3) (2002), 4303-4312.

[9] G. A. Garkusha, A. I. Generalov, Grothendieck categories as quotient categories of (Rmod,Ab), Fund. Appl. Math. 7(4) (2001), 983-992 (Russian).

[10] L. Gruson, C. U. Jensen, Dimensions cohomologiques reliées aux foncteurs lim^{(i)}, Séminaire d’algèbre (eds P. Dubreil and M.-P. Malliavin), Lecture Notes in Mathematics 867 (Springer, Berlin, 1981), pp. 234-294.

[11] I. Herzog, Elementary duality of modules, Trans. Amer. Math. Soc. 340 (1993), 37-69.

[12] I. Herzog, The Ziegler spectrum of a locally coherent Grothendieck category, Proc. London Math. Soc. 74(3) (1997), 503-558.

[13] M. Hovey, Model categories, American Mathematical Society, Providence, RI, 1999.

[14] M. Hovey, J. H. Palmieri, N. P. Strickland, Axiomatic stable homotopy theory, Mem. Amer. Math. Soc. 128 (1997), no. 610.

[15] L. Illusie, Existence de ré solutions globales, Sem. Geom. algebrique Bois Marie 1966/67, SGA VI, Lectre Notes in Mathematics 225, 1971, pp. 160-221.

[16] B. Keller, Derived categories and their uses, Handbook of algebra, Vol. 1, North-Holland, Amsterdam, 1996, pp. 671-701.

[17] M. Levine, Motivic cohomology and algebraic cycles a categorical approach, K-theory Preprint Archives 107, 1996.

[18] S. Mac Lane, Categories for the working mathematician, Graduate Texts in Mathematics 5, Springer-Verlag, New-York, 1971.

[19] A. Neeman, The Grothendieck duality theorem via Bousfield’s techniques and Brown representability, J. Amer. Math. Soc. 9(1) (1996), 205-236.

[20] M. Prest, Model theory and modules, London Math. Soc. Lecture Notes Series 130, Cambridge University Press, Cambridge, 1988.

[21] M. Prest, Purity, spectra and localisation, Encyclopedia of Mathematics and its Applications 121, Cambridge University Press, 2009.

[22] D. G. Quillen, Homotopical algebra, Lecture Notes in Mathematics 43, Springer-Verlag, Berlin-New York, 1967.

[23] S. Schwede, B. Shipley, Algebras and modules in monoidal model categories, Proc. London Math. Soc. 80(3) (2000), 491-511.

[24] B. Stenström, Rings of quotients, Grundlehren Math. Wiss., vol. 217, Springer-Verlag, New York–Heidelberg, 1975.

[25] C. A. Weibel, An introduction to homological algebra, Cambridge University Press, 1997.

Department of Mathematics, Swansea University, Singleton Park, Swansea SA2 8PP, United Kingdom

   E-mail address: hassan1167@yahoo.com

Department of Mathematics, Swansea University, Singleton Park, Swansea SA2 8PP, United Kingdom

   E-mail address: g.garkusha@swansea.ac.uk